OpenStudy (anonymous):

SOLVED! Find the matrix: A-3I (identity), where A= [4 -1] [7 -9]

OpenStudy (kainui):

They didn't, that's not the answer. A is a 2x2 matrix. There's no way you can add a scalar multiple of the identity matrix to get a 3x3 matrix. You must have something wrong.

OpenStudy (anonymous):

Yes,I did have something wrong. I thought I updated this, but I guess I should update all the sections. Thank you for your reply!

OpenStudy (kainui):

Yeah, glad you could figure it out.

OpenStudy (mathmate):

The "scalar" is actually 3I (not 31), or 3 times the identity matrix. The solution follows. \[\left[\begin{matrix}4 & -1 \\ 7 & -9\end{matrix}\right]-3\left[\begin{matrix}1 & 0 \\ 0 & 1\end{matrix}\right]=\left[\begin{matrix}4 & -1 \\ 7 & -9\end{matrix}\right]-\left[\begin{matrix}3 & 0 \\ 0 & 3\end{matrix}\right]=\left[\begin{matrix}1 & -1 \\ 7 & -12\end{matrix}\right]\]